Consider the differential equation X(1 — x)y + [p – (p+ 2)x]y – Py = 0, where p is a constant. (a) If p is not an integer,Write the general solution found in (a) in terms of elementary functions;

please be as neat as possible thank u
Answer #1

Sa avem deterrendal equation is 24(1-2); +[P-P+2 - Py = 0 la (2) To solve the given equ² , we first knows that, -2 (1-2) Yleze:(when d=1, p= P.D= Here 8=P is not an integer. The general sucution of (2) is y= autbre where U (8.0; 7; 2) = F(1, PP:)